(2x+3)dy/dx -y+(2x+3)
1
Expert's answer
2021-10-11T16:29:40-0400

"(2x+3)\frac{dy}{dx}-y=0\\
\Rightarrow \frac{dy}{dx}=\frac{y}{2x+3}\\
\Rightarrow \frac{dy}{y}=\frac{dx}{2x+3}"

Integrating both sides, we get:

"ln(y)=\frac{ln(2x+3)}{2}+ln(c)\\
\Rightarrow 2ln(y)=ln(2x+3)+2ln(c)\\
\Rightarrow ln(y^2)=ln(2x+3)+ln(c^2)\\
\Rightarrow ln(y^2)=ln(c^2(2x+3))\\
\therefore y^2=c^2(2x+3)\\
\Rightarrow y^2=C(2x+3)"